4) In a population of guppies living in a stream there is a color polymorphism. Homozygous tropical individuals have either blue or red stripes, depending on the allele they carry. Heterozygotes have sample 724 guppies from the pond. blue, and 126 are purple. purple stripes. You a) What is the frequency of the red allele in your sample? b) What is the frequency of the red genotype in your sample?
Answer #1

Answer:

a). The frequency of the red allele = fr(R) = 0.62

Explanation:

Explanation:

Red guppies (RR)= 724 – (216 + 126) = 382

Blue guppies (rr) = 216

Purple guppies (Rr) = 126

RR 328 = 382 * 2= 764 R alleles

rr 216 = 216 * 2 = 432 r alleles

Rr 126 = 126 R & 126 r alleles

Total alleles = 764 + 432 + 126 + 126 = 1448

Total R alleles = 764 + 126 = 890

Frequency of R allele = 890/1448 = 0.62

Total r alleles = 432+ 126 = 558

Frequency of r allele = 558/1448 = 0.38

b).

The expected frequency of the red genotype = fr (RR) = 0.62 * 0.62 = 0.38

The observed frequency of the red genotype = 382 / 724 = 0.53